# Memento 2: Learning by Stateful Reflective Memory

### v2

@@ 3.2.2 Transforming to Reflected MDP
It is worth mentioning that this augmented state admits a natural cognitive interpretation. The environment state $s_{t}$ encodes the agent’s external sensory or task-level signal, while the episodic memory $M_{t}$ represents an internal cognitive state that stores and organises past experiences, outcomes, and contextual information. The pair $(s_{t},M_{t})$ therefore constitutes the agent’s effective decision state, integrating external observation with internal experience. This perspective parallels belief-state constructions in partially observable Markov decision processes, where internal memory summarises relevant history to support optimal control [kaelbling1998planning], and aligns with cognitive theories that view episodic memory as a core substrate for human decision making and learning [tulving1983elements, GershmanDaw2017EpisodicRL].
@@ 4.1 Policy Iteration
We now define the retrieval mechanism that underpins the learning method. The central idea is to use local state density to guide which past cases should be retrieved from memory. To capture the similarity structure in the memory bank, we employ Parzen window estimation [Parzen1962, wang2008unified].
@@
where $\delta_{c_{\varnothing}}$ denotes the unit point mass at $c_{\varnothing}$. This mixture form shows that the agent automatically interpolates between reflection(retrieval)-driven behaviour when similarity to stored cases is high ($\lambda(x)\approx 1$) and knowledge-driven discovery through $c_{\varnothing}$ when similarity is low ($\lambda(x)\approx 0$). This mechanism mirrors the “discovery” component in infinite-armed bandits [berry1997bandit], complementing exploitation and exploration with the capacity to act beyond the current memory. This mechanism ensures that the agent can always act, even in novel states, while still exploiting relevant past experiences when available.
@@
$\mu_{k+1}(\cdot\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c}\nu(c\mid x)Q^{\mu_{t}}(x,c)-\alpha\mathrm{KL}(\nu(\cdot\mid x)\|\mu_{0}(\cdot\mid x))\right\},$
@@ 5 Related Work
The integration of memory mechanisms into reinforcement learning has long been explored to address the challenges of partial observability and sample complexity. Early approaches often relied on recurrent neural networks (RNNs), in particular Long Short-Term Memory (LSTM) networks [hochreiter1997long], to maintain hidden states that capture historical information. For instance, [hausknecht2015deep] extends DQN with an LSTM to summarise past observations for Q-learning under partial observability, providing an early neural-memory approach to POMDPs [kaelbling1998planning]. However, its memory is implicit and entangled in the recurrent hidden state, making long-horizon retention, interpretability, and revision of past information difficult.
More recent methods have taken inspiration from cognitive neuroscience, especially the hippocampal processes of memory formation and retrieval in humans and animals. For instance, the Hippocampal-Augmented Memory Integration (HAMI) framework [hami2025] incorporates symbolic indexing and hierarchical refinement based on hippocampal functions. Although HAMI shares our biological motivation, our Parzen-KL retrieval mechanism diverges in its probabilistic formulation and information-theoretic regularisation, which underpin both stability and scalability.
From another strand of memory-based RL, episodic control methods [ramani2019short] such as Model Free Episodic Control [blundell2016model] and Neural Episodic Control [pritzel2017neural] use memory primarily as a direct control mechanism, retrieving past transitions to approximate values or reuse actions via similarity matching. In this setting, memory effectively acts as a nonparametric policy, and generalisation is limited to local interpolation over stored experiences [ramani2019short].
@@
Recent work on Stable Hadamard Memory [hadamard2024] highlights the importance of stability in memory systems, using Hadamard products to calibrate and update stored information. While complementary in spirit, our method focuses instead on retrieval optimisation through information-theoretic regularisation. These two perspectives, stability of storage and optimality of retrieval, could be integrated in future research to build more comprehensive memory architectures.
@@
Case-based reasoning (CBR) [AamodtPlaza1994, Ashley1990] has long inspired memory-driven problem solving through the retrieval and adaptation of past cases. Our prior work on CBR-based LLMs [guo2024ds, guo2025optimizing] provides empirical evidence, particularly in data science and coding tasks, that LLMs are highly effective at revising retrieved cases, enabling strong generalisation to new tasks and overcoming a critical bottleneck in traditional CBR systems. Within our framework, CBR-based LLMs can be viewed as a stateless special case, for which we provide a principled and unified treatment.
@@
Experience replay, a widely used mechanism in deep RL since its adoption in DQN [mnih2015human], has evolved from uniform sampling to prioritised replay [schaul2015prioritized], where transitions are sampled according to temporal-difference errors. While these methods improve sample efficiency, they do not provide the structured, similarity-based retrieval required for generalisation across diverse contexts. Our approach diverges here by employing density-aware retrieval via Parzen window estimation, offering more nuanced and context-sensitive access to stored experiences.
@@
Our work also connects to broader developments in information-theoretic reinforcement learning. Maximum entropy RL [haarnoja2018soft] introduces entropy regularisation to encourage exploration and robustness, while KL constraints in algorithms such as TRPO [schulman2015trust] and MPO [abdolmaleki2018maximum] ensure stable policy updates. The distinctive aspect of our contribution is the application of KL regularisation not only to the policy but also to the retrieval process itself. This dual-level regularisation balances the exploitation of high-value memories with the exploration of diverse experiences, whilst preserving consistency with the underlying similarity structure.
@@
Finally, from a theoretical perspective, our framework extends established convergence results in reinforcement learning. The convergence proof for our soft policy iteration builds upon entropy-regularised MDPs [neu2017unified] but adapts them to the context of memory-based retrieval. Our two-time-scale analysis further draws on stochastic approximation theory [borkar2009stochastic], extending it to SRDPs where value and memory are updated simultaneously. A central outcome is the formal equivalence we establish between read (retrieval) and write operations and the fundamental processes of policy improvement and policy evaluation. This connection makes retrieval-augmented LLMs provably learnable for the first time, opening the way for reinforcement learning theory and tools to be systematically applied in this emerging setting.
@@ 6 Conclusions and Future Work
In this paper, we have established a formal theoretical foundation for continual learning in LLM agents through episodic memory and reflection, without requiring parameter fine-tuning. The core contribution is the Stateful Reflective Decision Process (SRDP), which extends the classical MDP framework by modeling reflection as a dual-action process: a retrieval stage that reads from episodic memory (policy improvement) and an action generation stage conditioned on retrieved cases. By augmenting the state space to include memory, we prove this system is Markovian and can be analysed as a Reflected MDP. We then introduce Read-Write Reflective Learning, a practical algorithm that integrates Parzen-window-based retrieval with KL-regularised soft policy iteration. Key theoretical results include: (1) guaranteed convergence of the policy iteration algorithm under bounded rewards, (2) two-timescale convergence when memory evolves slowly relative to policy updates, and (3) asymptotic optimality: as episodic memory grows to cover the state space, the composite policy provably converges to the optimal policy of the underlying MDP. This framework unifies case-based reasoning [guo2024ds], retrieval-augmented generation, and reinforcement learning under a single mathematical formalism, providing the first rigorous convergence guarantees for memory-driven LLM agents.
@@ 6.1 Internalising Episodic Memory into the LLM Architecture
In modern contexts, LLM agents increasingly employ explicit episodic memory systems to extend context and improve performance across diverse tasks [Zhong2024, Chhikara2025, fountas2024human]. Yet, the integration of memory retrieval with decision-making in LLMs still lacks a firm theoretical foundation, particularly regarding the conditions under which such systems converge to optimal behaviour. Our work addresses this gap directly.
@@
A natural next step is to integrate the episodic memory and read–write reflection mechanism of SRDP directly into the LLM architecture, yielding a single neural system whose internal state corresponds to the augmented state $(s,M)$. Recent “test-time memorisation” lines of work provide useful architectural blueprints. In particular, Titans introduces a neural long-term memory module that updates while the model is running, thereby supporting continual memorisation beyond the context window without offline retraining [behrouz2024titans]. MIRAS further reframes modern sequence models as instances of associative memory with explicit choices of memory structure, retention, and online optimisation [behrouz2025miras]. Our EM-LLM work in [fountas2024human] introduces a human-inspired episodic memory architecture that enables large language models to operate with effectively infinite context by selectively storing, retrieving, and reusing past experiences to support long-horizon reasoning and decision-making. These developments suggest a promising direction: replace external retrieval of SRDP over a growing case base by an internal memory state that is updated online, while retaining the SRDP control view in which “read” implements policy improvement and “write” implements policy evaluation.
@@
Our Read-Write loop differs from Titans-style test-time updates in a key way: SRDP performs stateful adaptation through experience-grounded reflection and memory evolution, rather than gradient-based parameter updates of the base model. A concrete research direction is to design a hybrid architecture in which the LLM remains fixed or slow-updated, but an internal memory module (or a small retrieval policy head) is updated online using environment feedback, with a provable two-time-scale separation between fast policy iteration and slow memory evolution. This would preserve the stability benefits of our reflected-MDP analysis while inheriting the efficiency and scalability motivations emphasised in Titans/MIRAS [behrouz2024titans, behrouz2025miras] and EM-LLM [fountas2024human].
@@
More broadly, recent work on test-time training layers and nested optimisation provides a conceptual bridge between architecture and learning dynamics. Test-time training (TTT) instantiates sequence layers whose hidden state is itself a learnable model updated during inference [sun2024ttt]. Nested Learning argues that modern deep models can be viewed as systems of nested optimisation problems with their own internal workflows, offering a principled lens on continual learning and catastrophic forgetting [behrouz2025nested]. An important future direction is to recast SRDP within this nested-optimisation view: the outer loop corresponds to fixed LLM representations and long-term objectives, while the inner loop implements read–write reflective adaptation as an online optimisation over an internal memory state. Such a unification could clarify when reflection-based memory updates are sufficient for continual learning, when parameter adaptation is necessary, and how to combine both without destabilising long-horizon control.
@@
Another dimension worth exploring is principled memory consolidation. Recent work on KV-state consolidation in Transformer architectures (Bottlenecked Transformers) demonstrates that selectively rewriting and compressing internal memory can improve predictive efficiency and long-horizon reasoning without expanding context length [oomerjee2025bottlenecked]. From the perspective of SRDP, such mechanisms provide an architectural realisation of the write operation, transforming episodic experience into a compact latent state that supports future reflection. This view is further supported by our work on Agent K, which shows that structured cycles of interaction, reflection, abstraction, and consolidation enable effective experiential learning without parameter fine-tuning [Grosnit2024KolbBasedEL]. Together, these results point toward an architecture in which episodic memory is consolidated on a slow time scale and queried on a fast time scale, aligning architectural dynamics with the SRDP formalism.
@@
Finally, these connections open new theoretical questions. MIRAS links sequence modelling to online optimisation and retention mechanisms [behrouz2025miras]; integrating SRDP into this framework may yield sharper characterisations of memory coverage, approximation error, and the conditions under which reflective memory updates guarantee monotone improvement and asymptotic optimality. Establishing such results for neural internal memories (rather than external case stores) would move SRDP toward a a unified neural state machine architecture in which episodic experience, reflection, and control are implemented end-to-end together.
@@ 7 Acknowledgements
We are grateful to Xue Yuan for valuable discussions on reinforcement learning and kernel methods, Siyuan Gu for insights into bandit algorithms and the discovery–exploration–exploitation trade-off, and Huichi Zhou and Yihang Chen and Yuxiang Chen for assistance with implementation and proof read. We also thank collaborators who reviewed earlier versions of this manuscript and provided helpful feedback.
@@ Appendix B Proof of Theorem 8
Unlike standard soft reinforcement learning [haarnoja2018soft], our setting involves state-dependent base measures that vary with memory content, requiring a careful analysis of the induced operators. In particular, extending KL-regularised evaluation requires establishing that the associated log-sum-exp operations preserve contraction for both the value function and the prior distributed resulted from memory updates. Lemma 16 extends the log-sum-exp Lipschitz bound [BoydVandenberghe2004] to this setting with Parzen base measures and memory updates.
@@ B.1 Bi-Lipschitz Properties
By the mean value theorem for vector-valued functions (e.g., [BoydVandenberghe2004]), there exists $\xi$ on the line segment between $q$ and $q^{\prime}$ such that
@@ B.3 Monotone Improvement of Read-Write Iterations
Our analysis relies on the Gibbs variational principle, also known as the log-sum-exp conjugate duality [BoydVandenberghe2004, Section 3.3], which connects entropy-regularised optimisation to exponential-family distributions. In particular, the Gibbs principle states that
@@
Our setting extends this principle by incorporating a base measure through the Kullback–Leibler divergence $\mathrm{KL}(\nu\|\mu_{0})$, replacing entropy regularisation with relative entropy. This yields a Parzen-style exponential-family formulation and corresponds to the Bregman divergence induced by negative entropy [Ziebart2010].

### v3

@@ 3.2.2 Transforming to Reflected MDP
It is worth mentioning that this augmented state admits a natural cognitive interpretation. The environment state $s_{t}$ encodes the agent’s external sensory or task-level signal, while the episodic memory $M_{t}$ represents an internal cognitive state that stores and organises past experiences, outcomes, and contextual information. The pair $(s_{t},M_{t})$ therefore constitutes the agent’s effective decision state, integrating external observation with internal experience. This perspective parallels belief-state constructions in partially observable Markov decision processes, where internal memory summarises relevant history to support optimal control [29], and aligns with cognitive theories that view episodic memory as a core substrate for human decision making and learning [50, 19].
@@ 4.1 Policy Iteration
We now define the retrieval mechanism that underpins the learning method. The central idea is to use local state density to guide which past cases should be retrieved from memory. To capture the similarity structure in the memory bank, we employ Parzen window estimation [38, 54].
@@
where $\delta_{c_{\varnothing}}$ denotes the unit point mass at $c_{\varnothing}$. This mixture form shows that the agent automatically interpolates between reflection(retrieval)-driven behaviour when similarity to stored cases is high ($\lambda(x)\approx 1$) and knowledge-driven discovery through $c_{\varnothing}$ when similarity is low ($\lambda(x)\approx 0$). This mechanism mirrors the “discovery” component in infinite-armed bandits [7], complementing exploitation and exploration with the capacity to act beyond the current memory. This mechanism ensures that the agent can always act, even in novel states, while still exploiting relevant past experiences when available.
@@
$\mu_{t+1}(\cdot\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c}\nu(c\mid x)Q^{\mu_{t}}(x,c)-\alpha\mathrm{KL}(\nu(\cdot\mid x)\|\mu_{0}(\cdot\mid x))\right\},$
@@ 5 Related Work
The integration of memory mechanisms into reinforcement learning has long been explored to address the challenges of partial observability and sample complexity. Early approaches often relied on recurrent neural networks (RNNs), in particular Long Short-Term Memory (LSTM) networks [27], to maintain hidden states that capture historical information. For instance, [26] extends DQN with an LSTM to summarise past observations for Q-learning under partial observability, providing an early neural-memory approach to POMDPs [29]. However, its memory is implicit and entangled in the recurrent hidden state, making long-horizon retention, interpretability, and revision of past information difficult.
More recent methods have taken inspiration from cognitive neuroscience, especially the hippocampal processes of memory formation and retrieval in humans and animals. For instance, the Hippocampal-Augmented Memory Integration (HAMI) framework [39] incorporates symbolic indexing and hierarchical refinement based on hippocampal functions. Although HAMI shares our biological motivation, our Parzen-KL retrieval mechanism diverges in its probabilistic formulation and information-theoretic regularisation, which underpin both stability and scalability.
From another strand of memory-based RL, episodic control methods [42] such as Model Free Episodic Control [9] and Neural Episodic Control [40] use memory primarily as a direct control mechanism, retrieving past transitions to approximate values or reuse actions via similarity matching. In this setting, memory effectively acts as a nonparametric policy, and generalisation is limited to local interpolation over stored experiences [42].
@@
Recent work on Stable Hadamard Memory [30] highlights the importance of stability in memory systems, using Hadamard products to calibrate and update stored information. While complementary in spirit, our method focuses instead on retrieval optimisation through information-theoretic regularisation. These two perspectives, stability of storage and optimality of retrieval, could be integrated in future research to build more comprehensive memory architectures.
@@
Case-based reasoning (CBR) [1, 3] has long inspired memory-driven problem solving through the retrieval and adaptation of past cases. Our prior work on CBR-based LLMs [23, 24] provides empirical evidence, particularly in data science and coding tasks, that LLMs are highly effective at revising retrieved cases, enabling strong generalisation to new tasks and overcoming a critical bottleneck in traditional CBR systems. Within our framework, CBR-based LLMs can be viewed as a stateless special case, for which we provide a principled and unified treatment.
@@
Experience replay, a widely used mechanism in deep RL since its adoption in DQN [34], has evolved from uniform sampling to prioritised replay [44], where transitions are sampled according to temporal-difference errors. While these methods improve sample efficiency, they do not provide the structured, similarity-based retrieval required for generalisation across diverse contexts. Our approach diverges here by employing density-aware retrieval via Parzen window estimation, offering more nuanced and context-sensitive access to stored experiences.
@@
Our work also connects to broader developments in information-theoretic reinforcement learning. Maximum entropy RL [25] introduces entropy regularisation to encourage exploration and robustness, while KL constraints in algorithms such as TRPO [45] and MPO [2] ensure stable policy updates. The distinctive aspect of our contribution is the application of KL regularisation not only to the policy but also to the retrieval process itself. This dual-level regularisation balances the exploitation of high-value memories with the exploration of diverse experiences, whilst preserving consistency with the underlying similarity structure.
@@
Finally, from a theoretical perspective, our framework extends established convergence results in reinforcement learning. The convergence proof for our soft policy iteration builds upon entropy-regularised MDPs [35] but adapts them to the context of memory-based retrieval. Our two-time-scale analysis further draws on stochastic approximation theory [12], extending it to SRDPs where value and memory are updated simultaneously. A central outcome is the formal equivalence we establish between read (retrieval) and write operations and the fundamental processes of policy improvement and policy evaluation. This connection makes retrieval-augmented LLMs provably learnable for the first time, opening the way for reinforcement learning theory and tools to be systematically applied in this emerging setting.
@@ 6 Conclusions and Future Work
In this paper, we have established a formal theoretical foundation for continual learning in LLM agents through episodic memory and reflection, without requiring parameter fine-tuning. The core contribution is the Stateful Reflective Decision Process (SRDP), which extends the classical MDP framework by modeling reflection as a dual-action process: a retrieval stage that reads from episodic memory (policy improvement) and an action generation stage conditioned on retrieved cases. By augmenting the state space to include memory, we prove this system is Markovian and can be analysed as a Reflected MDP. We then introduce Read-Write Reflective Learning, a practical algorithm that integrates Parzen-window-based retrieval with KL-regularised soft policy iteration. Key theoretical results include: (1) guaranteed convergence of the policy iteration algorithm under bounded rewards, (2) two-timescale convergence when memory evolves slowly relative to policy updates, and (3) asymptotic optimality: as episodic memory grows to cover the state space, the composite policy provably converges to the optimal policy of the underlying MDP. This framework unifies case-based reasoning [23], retrieval-augmented generation, and reinforcement learning under a single mathematical formalism, providing the first rigorous convergence guarantees for memory-driven LLM agents.
@@ 6.1 Internalising Episodic Memory into the LLM Architecture
In modern contexts, LLM agents increasingly employ explicit episodic memory systems to extend context and improve performance across diverse tasks [63, 15, 18]. Yet, the integration of memory retrieval with decision-making in LLMs still lacks a firm theoretical foundation, particularly regarding the conditions under which such systems converge to optimal behaviour. Our work addresses this gap directly.
@@
A natural next step is to integrate the episodic memory and read–write reflection mechanism of SRDP directly into the LLM architecture, yielding a single neural system whose internal state corresponds to the augmented state $(s,M)$. Recent “test-time memorisation” lines of work provide useful architectural blueprints. In particular, Titans introduces a neural long-term memory module that updates while the model is running, thereby supporting continual memorisation beyond the context window without offline retraining [6]. MIRAS further reframes modern sequence models as instances of associative memory with explicit choices of memory structure, retention, and online optimisation [4]. Our EM-LLM work in [18] introduces a human-inspired episodic memory architecture that enables large language models to operate with effectively infinite context by selectively storing, retrieving, and reusing past experiences to support long-horizon reasoning and decision-making. These developments suggest a promising direction: replace external retrieval of SRDP over a growing case base by an internal memory state that is updated online, while retaining the SRDP control view in which “read” implements policy improvement and “write” implements policy evaluation.
@@
Our Read-Write loop differs from Titans-style test-time updates in a key way: SRDP performs stateful adaptation through experience-grounded reflection and memory evolution, rather than gradient-based parameter updates of the base model. A concrete research direction is to design a hybrid architecture in which the LLM remains fixed or slow-updated, but an internal memory module (or a small retrieval policy head) is updated online using environment feedback, with a provable two-time-scale separation between fast policy iteration and slow memory evolution. This would preserve the stability benefits of our reflected-MDP analysis while inheriting the efficiency and scalability motivations emphasised in Titans/MIRAS [6, 4] and EM-LLM [18].
@@
More broadly, recent work on test-time training layers and nested optimisation provides a conceptual bridge between architecture and learning dynamics. Test-time training (TTT) instantiates sequence layers whose hidden state is itself a learnable model updated during inference [48]. Nested Learning argues that modern deep models can be viewed as systems of nested optimisation problems with their own internal workflows, offering a principled lens on continual learning and catastrophic forgetting [5]. An important future direction is to recast SRDP within this nested-optimisation view: the outer loop corresponds to fixed LLM representations and long-term objectives, while the inner loop implements read–write reflective adaptation as an online optimisation over an internal memory state. Such a unification could clarify when reflection-based memory updates are sufficient for continual learning, when parameter adaptation is necessary, and how to combine both without destabilising long-horizon control.
@@
Another dimension worth exploring is principled memory consolidation. Recent work on KV-state consolidation in Transformer architectures (Bottlenecked Transformers) demonstrates that selectively rewriting and compressing internal memory can improve predictive efficiency and long-horizon reasoning without expanding context length [36]. From the perspective of SRDP, such mechanisms provide an architectural realisation of the write operation, transforming episodic experience into a compact latent state that supports future reflection. This view is further supported by our work on Agent K, which shows that structured cycles of interaction, reflection, abstraction, and consolidation enable effective experiential learning without parameter fine-tuning [21]. Together, these results point toward an architecture in which episodic memory is consolidated on a slow time scale and queried on a fast time scale, aligning architectural dynamics with the SRDP formalism.
@@
Finally, these connections open new theoretical questions. MIRAS links sequence modelling to online optimisation and retention mechanisms [4]; integrating SRDP into this framework may yield sharper characterisations of memory coverage, approximation error, and the conditions under which reflective memory updates guarantee monotone improvement and asymptotic optimality. Establishing such results for neural internal memories (rather than external case stores) would move SRDP toward a unified neural state machine architecture in which episodic experience, reflection, and control are implemented end-to-end together.
@@ 7 Acknowledgements
We are grateful to Xue Yan for valuable discussions on reinforcement learning and kernel methods, Siyuan Guo for insights into bandit algorithms and the discovery–exploration–exploitation trade-off, and Huichi Zhou and Yihang Chen and Yuxiang Chen for assistance with implementation and proof-read. We also thank collaborators who reviewed earlier versions of this manuscript and provided helpful feedback.
@@ Appendix B Proof of Theorem 8
Unlike standard soft reinforcement learning [25], our setting involves state-dependent base measures that vary with memory content, requiring a careful analysis of the induced operators. In particular, extending KL-regularised evaluation requires establishing that the associated log-sum-exp operations preserve contraction for both the value function and the prior distributed resulted from memory updates. Lemma 16 extends the log-sum-exp Lipschitz bound [13] to this setting with Parzen base measures and memory updates.
@@ B.1 Bi-Lipschitz Properties
By the mean value theorem for vector-valued functions (e.g., [13]), there exists $\xi$ on the line segment between $q$ and $q^{\prime}$ such that
@@ B.3 Monotone Improvement of Read-Write Iterations
Our analysis relies on the Gibbs variational principle, also known as the log-sum-exp conjugate duality [13, Section 3.3], which connects entropy-regularised optimisation to exponential-family distributions. In particular, the Gibbs principle states that
@@
Our setting extends this principle by incorporating a base measure through the Kullback–Leibler divergence $\mathrm{KL}(\nu\|\mu_{0})$, replacing entropy regularisation with relative entropy. This yields a Parzen-style exponential-family formulation and corresponds to the Bregman divergence induced by negative entropy [65].